OpenStudy (anonymous):

if i have cos(4*pi*x - n*pi*x) what property or rule allows me to rewrite it as cos(4-n)*pi*x ?

OpenStudy (anonymous):

how about factoring the \(\pi x\) ?

OpenStudy (anonymous):

um okay. what rule states you can factor something out of a trig function like that?

OpenStudy (anonymous):

not the function.. just the angle in the function \[ \cos\left(4\pi x-n\pi x\right)=\cos\left[(4-n)\pi x\right] \]

OpenStudy (anonymous):

okay how does that angle end up on the otherside like \[\cos(4-n)\pi*x\]

OpenStudy (anonymous):

better question, help me integrate : \[\int\limits_{-1}^{0}\cos(4 \pi x)\cos(n \pi x) dx\]

OpenStudy (anonymous):

use the addition rule \[\cos C\cdot\cos D={1\over2}\left[\cos(C+D)+\cos(C-D)\right]\]